OpenStudy (maryfitzgerald14):

What can you add to 70 to make 90?

Nnesha (nnesha):

do opposite way what is 90 - 70 = ??

OpenStudy (quickstudent):

You can find out simply by subtracting :90 - 70 = ?

OpenStudy (maryfitzgerald14):

isnt the answer 20

Nnesha (nnesha):

okay now check your answer 70 + 20 = ????

OpenStudy (maryfitzgerald14):

is it 90

Nnesha (nnesha):

yes so that's right :)

OpenStudy (maryfitzgerald14):

ok thanks

OpenStudy (quickstudent):

correct:)